# Closed-Form Interpretation of Neural Network Latent Spaces with Symbolic Gradients

### v1

@@ Abstract
It has been demonstrated in many scientific fields that artificial neural networks like autoencoders or Siamese networks encode meaningful concepts in their latent spaces. However, there does not exist a comprehensive framework for retrieving this information in a human-readable form without prior knowledge. In order to extract these concepts, we introduce a framework for finding closed-form interpretations of neurons in latent spaces of artificial neural networks. The interpretation framework is based on embedding trained neural networks into an equivalence class of functions that encode the same concept. We interpret these neural networks by finding an intersection between the equivalence class and human-readable equations defined by a symbolic search space. The approach is demonstrated by retrieving invariants of matrices and conserved quantities of dynamical systems from latent spaces of Siamese neural networks.
@@ 2 Related Work
Our article aims to interpret existing latent space models. We extend an interpretation framework Wetzel (2024) originally developed to interpret neural network classifiers to interpret neural network latent spaces.
@@ 3.1 Siamese Neural Networks
$\mathcal{L}(A,P,N)=\max(d(f(A),f(P))-d(f(A),f(N))+\alpha,0).$
The anchor $A$ is the ground truth class, the positive sample $P$ is of the same class as $A$, whereas the negative sample $N$ is of a different class. Instead of using a twin network, this setup requires a triplet of identical networks, each still sharing the same weights. The triplet loss is minimized when the distance between the anchor and positive sample is minimized in the embedding space, while the distance between the anchor and negative sample is maximized. The margin parameter $\alpha$ is a positive constant which encourages separation between positive and negative samples, as $\alpha=0$ would mean that the loss could be trivially minimized by projecting all samples to the same location. Finally, the $\max(\cdot)$ operation ensures that the distance between positive and negative samples remains finite.
@@ 4.1.3 Invariants Under the Lorentz Transformation
In experiment 6, we apply the Lorentz transformation to the field strength tensor $F_{\mu\nu}$, which gives rise to the Lorentz invariants X and Y. Since the antisymmetry of $F_{\mu\nu}$ is preserved under the Lorentz transformation, each member of a triplet is antisymmetric, so we only use the 6 off-diagonal entries above (or equivalently below) the main diagonal as our input to the neural network. The anchor is a vector of these 6 entries from $F_{\mu\nu}$,

### v2

@@ Abstract
It has been demonstrated in many scientific fields that artificial neural networks, like autoencoders or Siamese networks, encode meaningful concepts in their latent spaces. However, there does not exist a comprehensive framework for retrieving this information in a human-readable form without prior knowledge. In order to extract these concepts, we introduce a framework for finding closed-form interpretations of neurons in latent spaces of artificial neural networks. The interpretation framework is based on embedding trained neural networks into an equivalence class of functions that encode the same concept. We interpret these neural networks by finding an intersection between the equivalence class and human-readable equations defined by a symbolic search space. The approach is demonstrated by retrieving invariants of matrices and conserved quantities of dynamical systems from latent spaces of Siamese neural networks.
@@ 2 Related Work
Our article aims to interpret existing latent space models. We extend an interpretation framework Wetzel (2024), originally developed to interpret neural network classifiers, to interpret neural network latent spaces.
@@ 3.1 Siamese Neural Networks
$\mathcal{L}(x_{A},x_{P},x_{N})=\max(d(f(x_{A}),f(x_{P}))-d(f(x_{A}),f(x_{N}))+\alpha,0).$
The anchor $x_{A}$ is the ground truth class, the positive sample $x_{P}$ is of the same class as $x_{A}$, whereas the negative sample $N$ is of a different class. Instead of using a twin network, this setup requires a triplet of identical networks, each still sharing the same weights. The triplet loss is minimized when the distance between the anchor and positive sample is minimized in the embedding space, while the distance between the anchor and negative sample is maximized. The margin parameter $\alpha$ is a positive constant which encourages separation between positive and negative samples, as $\alpha=0$ would mean that the loss could be trivially minimized by projecting all samples to the same location. Finally, the $\max(\cdot)$ operation ensures that the distance between positive and negative samples remains finite.
@@ 4.1.3 Invariants Under the Lorentz Transformation
In experiment 6, we apply the Lorentz transformation to the field strength tensor $F_{\mu\nu}$, which gives rise to the Lorentz invariants in Appendix A.1.3. Since the antisymmetry of $F_{\mu\nu}$ is preserved under the Lorentz transformation, each member of a triplet is antisymmetric, so we only use the 6 off-diagonal entries above (or equivalently below) the main diagonal as our input to the neural network. The anchor is a vector of these 6 entries from $F_{\mu\nu}$,